# Capturing reduced-order quantum many-body dynamics out of equilibrium via neural ordinary differential equations

### v1

@@ 2.2.1 Cumulant expansion
Any ppRDM can be decomposed into antisymmetrised products of lower‑order RDMs plus a connected part (the cumulant) [kutzelnigg1999Cumulant]. For the two-particle reduced density matrix the decomposition reads
@@
where $\hat{A}$ antisymmetrises the product and $\Delta^{(2)}_{12}$ is the two‑particle cumulant. For the three-particle reduced density matrix one obtains
@@
where $D_{123}$ is the three‑particle cumulant (the genuinely connected three‑body correlation). We have chosen a convention for $\hat{A}$ such that every topologically distinct term contributes only ones in order to avoid normalization in the cumulant expansion Eq. (6). The first term contains only products of the 1RDM so that the three particles are uncorrelated in their dynamics. The second term shows that once the two‑particle cumulant $\Delta_{12}$ is known, all pairwise correlations in the three‑body system are already determined; only the last term $\Delta_{123}$ adds the correlated three‑body part. As has been shown in Ref. [lackner2015Propagating], parts of the three-particle cumulant $\Delta_{123}$ can be exactly reconstructed from the two-particle reduced density matrix through the knowledge of its traces. The irreducible three-body part of $\Delta_{123}$ denoted as $\Delta_{123,\text{K}}$ with $\text{Tr}_{3}\Delta_{123,\text{K}}=0$, however, cannot be expressed through lower‑order RDMs. Consequently, an accurate reconstruction of $D_{123}$ depends on a good approximation of $\Delta_{123,\text{K}}$, while the bulk of the many‑body information is already encoded in $D_{12}$ (through $\Delta_{12}$).
All practical time-dependent two-particle reduced density matrix implementations close the Bogoliubov–Born–Green–Kirkwood–Yvon hierarchy by replacing the three-particle reduced density matrix $D_{123}$ with a functional of the two-particle reduced density matrix $D_{12}$ (i.e. of the two-particle cumulant $\Delta_{12}$). The most widely employed approximations are time‑local, thereby neglecting any memory dependence. A more detailed derivation and specific functional forms can be found in Ref. [donsa2023Nonequilibrium].
2.3 Neural ordinary differential equation
Neural ordinary differential equations provide a data‑driven framework to learn a continuous‑time dynamical system from sampled trajectories [chen2018Neural]. Given a set of discrete two-particle reduced density matrix snapshots $\{D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}(t_{k})\}_{k=0}^{K}$ obtained from the propagation of the wavefunction through exact diagonalization of the Schrödinger equation, a neural ordinary differential equation seeks a parametric vector field $\mathcal{F}_{\theta}\bigl(D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}\bigr)$ such that
@@ 2.3 Neural ordinary differential equation
with $\Phi_{\theta}$ the solution of the neural ODE obtained by an adaptive ODE solver and $\|\cdot\|_{F}$ the Frobenius (Schatten‑2) norm. After training, the learned vector field can be evaluated on unseen initial conditions to test whether the dynamics is well described by a time‑local functional of the instantaneous two-particle reduced density matrix.
Thus, neural ordinary differential equations serve as an independent, model‑agnostic diagnostic tool to map the parameter regimes (in the UUVV plane) where the time-dependent two-particle reduced density matrix closure can be reliably approximated by time‑local reconstructions as discussed in Sec. 2.2.1.
In practice, we employ a fully connected architecture with two hidden layers, each with a dimension of $2048$. The mixed-spin block of the two-particle reduced density matrix $D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}$ contains $1296$ complex-valued entries. To pass them through the neural ordinary differential equation, we must separate the real and imaginary parts, yielding a total of $2592$ real values. This can be reduced by leveraging its hermeticity and taking only the upper triangular values, which in turn results in $1296$ real values. For each parameter configuration in the $U$-$V$ plane, we divide our time series into training, validation, and test sets. The first $3000$ time steps are used for training and the next $1000$ time steps are used for validation, where one time step is $\Delta t=0.01\,J^{-1}$. The remainder can be used for testing. We apply min-max normalization to the real and imaginary parts separately to scale them to $[0,1]$. Here, we first normalize the train set, then use the same normalization constants for both the validation and test sets. We use the Python library torchdiffeq [torchdiffeq] for the implementation of the neural ordinary differential equation.
It is important to mention that, compared to other applications of neural ordinary differential equations, which typically reduce high-dimensional data to a lower-dimensional manifold and then train the neural ordinary differential equation on the low-dimensional representation [guo2025Efficient, lai2022Neural, norcliffe2021Neural, sholokhov2023Physicsinformed, lee2021Parameterized], we train the model directly on the high-dimensional data. Furthermore, the size of the training set is of the same order of magnitude as the dimensionality. To our best knowledge, this is the first time that neural ordinary differential equations has been trained in this way, which, if successful, could open up new application possibilities.
@@ 3.1 Time-local reconstruction for different parameter configurations
To determine where a time‑local reconstruction of the three‑particle cumulant $\Delta_{123}(t)$ is justified, we first analyze three quantitative indicators obtained from the exact dynamics of the two‑ and the kernel component of the three‑particle cumulants, $\Delta_{12}(t)$ and $\Delta_{123,\text{K}}(t)$, for a broad range of interaction strengths $U$ and quench amplitudes $V$. These measures serve both to interpret the neural ordinary differential equation results and to determine the general structure of the $U$-$V$ parameter plane.
@@
$\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=\frac{1}{T}\int_{0}^{T}\left\|\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}(t)\right\|-\left\|\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}(0)\right\|\text{d}t\,$, (9)
where $\|\cdot\|$ denotes the Frobenius norm. The contour $\delta\Delta_{123^{\prime}\text{K}}=0.65$ separates regimes of moderate ($\delta\Delta_{123,\text{K}}\leq 0.65$) and strong ($\delta\Delta_{123,\text{K}}>0.65$) three‑particle correlations buildup, as visible in Fig. 2(a). This choice is empirical as it delimits the regimes very well. All three quantities are evaluated for $T=50\,J^{-1}$ for the whole parameter scan.
The instantaneous two‑particle correlation energy is defined as
@@
in Fig. 2(b). The colour scale is truncated at 0.1-0.1 for visual clarity; values below this threshold are present but not displayed. Throughout the explored parameter space, the correlation energy is initially negative and increases during the time evolution, attaining positive average values for some parameter configurations. Figure 2(b) thus visualizes the buildup of two‑particle correlations. The cut‑off at 0.1-0.1 was chosen empirically to separate regimes of pronounced correlation‑energy increase from those exhibiting only a modest rise. Both contour lines in Fig. 2 indicate structures in the dynamics of the two-particle reduced density matrix across the parameter plane, which serve as reference lines for better comparability across plots.
@@
Figure 2: We study the six-site Fermi-Hubbard model of Eq. (1). Both plots show a parameter scan over the initial potential strength $V$ and the interaction $U$. The dynamical buildup of three-particle correlations as described in Eq. (9) is visualized in (a). The solid black line corresponds to $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$, which indicates the crossover between moderate and strong correlation buildup. In (b) the time-averaged correlation energy relative to the initial excitation energy $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)$ is shown, where the black dashed line corresponds to $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$. This delimits the regime of strong two-particle correlation buildup. Both values were chosen in accordance with Fig. 3 since they also delimit the positive and negative as well as the strong and weak correlations of the cumulant dynamics, respectively.
@@
We hypothesize that in regions with high (anti-)correlation between two-particle and three-particle cumulants, the neural ordinary differential equation should perform well: The information about the three-particle cumulant, which is required to close the equations of motion for the two-particle reduced density matrix, should be contained in the two-particle reduced density matrix. In the anti-correlated regime, there is a strong correlation between these two quantities, though negative. This could hint at the existence of a time-local reconstruction functional in the anti-correlated regime since the current reconstructions only perform well in the positively correlated regime. The existence of such a functional would allow efficient numerical integration of the equations of motion with accurate results in this regime. The neural ordinary differential equation serves as a model-agnostic tool: It can only learn Markovian dynamics by design. Thus, if memory is required to reconstruct the three-particle cumulant and hence close the equations of motion, the model should not be able to learn and predict the exact dynamics of the two-particle reduced density matrix.
@@
To assess this hypothesis, we look at the Pearson correlation coefficient $C_{f,g}$ [pearson1896Mathematical], which provides a quantitative measure of the linear relationship between two time‑dependent observables $f(t)$ and $g(t)$. It is defined as the normalized covariance
@@
By construction, $1\leq C_{f,g}\leq 1$; $C_{f,g}=1$ (1-1) indicates perfect (anti‑)correlation, while $C_{f,g}=0$ corresponds to the absence of linear correlation. In the present work, $C_{f,g}$ is employed to assess the temporal correlation between the two‑particle cumulant $\Delta_{12}(t)$ and the kernel component of the three‑particle cumulant $\Delta_{123,\text{K}}(t)$. Since the initial rise of the cumulants is always correlated and we are only interested in the correlations of the steady-state fluctuations, we omit the start of the time series and empirically choose $t_{0}=10\,J^{-1}$. According to our hypothesis, a high absolute value of $C_{\Delta_{12},\Delta_{123,\text{K}}}$ supports the validity of time‑local reconstruction functionals, whereas low absolute values signal the need for memory‑dependent extensions. In Fig. 3(a) and (b) the Pearson correlation coefficient between the kernel component of the three-particle cumulant $\Delta_{123,\text{K}}^{\uparrow\uparrow\downarrow}$ and the two-particle cumulants $\Delta_{12}^{\uparrow\uparrow}$ and $\Delta_{12}^{\uparrow\downarrow}$, respectively, are shown. Both contour lines, which are equal to those in Fig. 2, indicate the crossover between positive and negative correlation ($\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$), and to the region of strong time correlation ($\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$). Note that evaluating in Fig. 2 and 3 the entire three-particle cumulant instead of only the kernel component (such as in Ref. [donsa2023Nonequilibrium]) leads to almost the same results because the trace information is a small, practically time-independent contribution.
Figure 3: In (a) and (b) the Pearson correlation coefficient $C_{\Delta_{12},\Delta_{123,\text{K}}}$ between the kernel component of the three-particle cumulant $\Delta_{123,\text{K}}^{\uparrow\uparrow\downarrow}$ and the two-particle cumulants $\Delta_{12}^{\uparrow\uparrow}$ and $\Delta_{12}^{\uparrow\downarrow}$ are shown, respectively, across the $U$-$V$ plane for the six-site Fermi-Hubbard model of Eq. (1). The contour lines $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$ and $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$ from Fig. 2 are visualized as well. In (c) and (d) the Pearson correlation coefficient between the predicted and the exact time series of the two-particle reduced density matrix $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ is visualized for a prediction length of $t_{\text{pred}}=20\,J^{-1}$ and $t_{\text{pred}}=25\,J^{-1}$ (2000 and 2500 time steps), respectively, starting from $D_{12}(t=40\,J^{-1})$. For each parameter configuration, one model was trained.
We use $D_{12}(t_{0}=40\,J^{-1})$ as the initial input to the neural ordinary differential equation, which is then employed to predict the subsequent evolution of the two-particle reduced density matrix starting from $t_{0}=40\,J^{-1}$. Consequently, we define the prediction length as
@@
To quantify the predictive performance of the neural ordinary differential equation, we calculate the Pearson correlation coefficient between the predicted time series of the two-particle reduced density matrix and the actual time series for the reduced representation (upper triangular values only) $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ for a prediction length of $t_{\text{pred}}=20\,J^{-1}$ and $t_{\text{pred}}=25\,J^{-1}$ (2000 and 2500 time steps) starting from $D_{12}(t=40\,J^{-1})$, visualized in Fig. 3(c) and (d), respectively. The contour lines are the same as in Fig. 2. For a prediction length of $t_{\text{pred}}=20\,J^{-1}$ the performance is high in the correlated regime and reasonably good in the anti-correlated regime. However, for a longer prediction length of $t_{\text{pred}}=25\,J^{-1}$, the performance in the anti-correlated regime drops significantly, while it remains good in the correlated regime.
To further investigate the accuracy of the prediction, we calculate two quantities from the predicted time series, namely the occupation number $n(t)$ and the doublon occupation number $d(t)$ (number of pairs on a site). Since all sites show similar dynamics, we only consider site 1. The accuracy of the prediction is determined by the time-integrated deviations from the exact result
@@
This quantity as well as the deviations for the doublon occupation number $\overline{\delta d_{1}}$ are shown in Fig. 4 for both prediction lengths. Again, we can see that the error is low in the correlated regime, while it is high in the anti-correlated regime. Figure 4 also depicts that even for a shorter prediction length, there is a significant difference in the accuracy of the prediction in the correlated and anti-correlated regimes.
These results contradict our hypothesis and thus suggest that there might be no time-local reconstruction functional in the anti-correlated regime. The solid contour line separates moderate from strong correlations buildup of the kernel component of the three-particle cumulant ($\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$). In both Fig. 3 and 4, this line quite accurately localizes regions of good and bad performance, indicating that the existence of a time-local reconstruction functional depends on this quantity.

### v2

@@ 2.2.1 Cumulant expansion
Any ppRDM can be decomposed into antisymmetrised products of lower‑order RDMs plus a connected part (the cumulant) [54]. For the 2RDM the decomposition reads
@@
where $\hat{A}$ antisymmetrises the product and $\Delta^{(2)}_{12}$ is the two‑particle cumulant. For the 3RDM one obtains
@@
where $\Delta_{123}$ is the three‑particle cumulant (the genuinely connected three‑body correlation). We have chosen a convention for $\hat{A}$ such that every topologically distinct term contributes only ones in order to avoid normalization in the cumulant expansion Eq. (7). The first term contains only products of the 1RDM so that the three particles are uncorrelated in their dynamics. The second term shows that once the two‑particle cumulant $\Delta_{12}$ is known, all pairwise correlations in the three‑body system are already determined; only the last term $\Delta_{123}$ adds the correlated three‑body part. As has been shown in Ref. [55], parts of the three-particle cumulant $\Delta_{123}$ can be exactly reconstructed from the 2RDM through the knowledge of its traces. The irreducible three-body part of $\Delta_{123}$ denoted as $\Delta_{123,\text{K}}$ with $\text{Tr}_{3}\Delta_{123,\text{K}}=0$, however, cannot be expressed through lower‑order RDMs. Consequently, an accurate reconstruction of $D_{123}$ depends on a good approximation of $\Delta_{123,\text{K}}$, while the bulk of the many‑body information is already encoded in $D_{12}$ (through $\Delta_{12}$).
All practical TD2RDM implementations close the BBGKY hierarchy by replacing the 3RDM $D_{123}$ with a functional of the 2RDM $D_{12}$ (i.e. of the two-particle cumulant $\Delta_{12}$). All presently employed approximations are time‑local, thereby neglecting any memory dependence. A more detailed derivation and specific functional forms can be found in Ref. [34].
2.3 Neural ODE
Neural ODEs provide a data‑driven framework to learn a continuous‑time dynamical system from sampled trajectories [24]. Given a set of discrete 2RDM snapshots $\{D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}(t_{k})\}_{k=0}^{K}$ obtained from the propagation of the wavefunction through exact diagonalization of the Schrödinger equation, a neural ODE seeks a parametric vector field $\mathcal{F}_{\theta}\bigl(D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}\bigr)$ such that
@@ 2.3 Neural ODE
with $\Phi_{\theta}$ the solution of the neural ODE obtained by an adaptive ODE solver and $\|\cdot\|_{F}$ the Frobenius (Schatten‑2) norm. After training, the learned vector field can be evaluated on unseen initial conditions to test whether the dynamics is well described by a time‑local functional of the instantaneous 2RDM.
Thus, neural ODEs serve as an independent, model‑agnostic diagnostic tool to map the parameter regimes (in the UUVV plane) where the TD2RDM closure can be reliably approximated by time‑local reconstructions as discussed in Sec. 2.2.1.
In practice, we employ a fully connected architecture with two hidden layers, each with a dimension of $2048$. We tested deeper and wider feed-forward neural ODE variants, but did not observe meaningful differences in predictive behavior, hence we stick to this architecture. The mixed-spin block of the 2RDM $D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}$ contains $1296$ complex-valued entries. To pass them through the neural ODE, we must separate the real and imaginary parts, yielding a total of $2592$ real values. This can be reduced by leveraging its hermiticity and taking only the upper triangular values, which in turn results in $1296$ real values. For each parameter configuration in the $U$-$V$ plane, we divide our time series into training, validation, and test sets. The first $3000$ time steps are used for training and the next $1000$ time steps are used for validation, where one time step is $\Delta t=0.01\,J^{-1}$. The remainder can be used for testing. We apply min-max normalization to the real and imaginary parts separately to scale them to $[0,1]$. Here, we first normalize the train set, then use the same normalization constants for both the validation and test sets. We use the Python library torchdiffeq [25] for the implementation of the neural ODE.
It is important to mention that, compared to other applications of neural ODEs, which typically reduce high-dimensional data to a lower-dimensional manifold and then train the neural ODE on the low-dimensional representation [41, 57, 70, 90, 59], we train the model directly on the high-dimensional data. We intentionally do not reduce the dimensionality of the data, since we want to test the Markovianity of the dynamics. If we were to reduce the dimensionality, and then train the neural ODE on the low-dimensional representation, we would not be able to distinguish whether a failure of the neural ODE is due to a non-Markovian dynamics or due to information loss in the dimensionality reduction, resulting in a non-Markovian dynamics in the low-dimensional representation. By training the neural ODE directly on the high-dimensional data, we can directly test the Markovianity of the dynamics without this ambiguity. Furthermore, the size of the training set is of the same order of magnitude as the dimensionality. To our best knowledge, this is the first time that neural ODEs have been trained in this way, which, if successful, could open up new application possibilities.
@@ 3.1 Time-local reconstruction for different parameter configurations
To determine where a time‑local reconstruction of the three‑particle cumulant $\Delta_{123}(t)$ is justified, we first analyze three quantitative indicators obtained from the exact dynamics of the two‑particle cumulant, $\Delta_{12}(t)$, and the kernel component of the three‑particle cumulants, $\Delta_{123,\text{K}}(t)$, for a broad range of interaction strengths $U$ and quench amplitudes $V$. These measures serve both to interpret the neural ODE results and to determine the general structure of the $U$-$V$ parameter plane.
@@
$\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=\frac{1}{T}\int_{0}^{T}\left\|\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}(t)\right\|_{F}-\left\|\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}(0)\right\|_{F}\text{d}t\,$, (10)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. The contour $\delta\Delta_{123,\text{K}}=0.65$, which can be used to separate regimes of moderate ($\delta\Delta_{123,\text{K}}\leq 0.65$) and strong ($\delta\Delta_{123,\text{K}}>0.65$) three‑particle correlations buildup, as visible in Fig. 2(a), can be simultaneously used to delimit regimes of positive and negative correlations in the dynamics of the cumulants, see Fig. 3. For the upper time limit we take $T=50\,J^{-1}$ for the whole parameter scan.
As another measure of correlations buildup, we use the instantaneous two‑particle correlation energy defined as
@@
in Fig. 2(b). The color scale is truncated at 0.1-0.1 for visual clarity; values below this threshold are present but not displayed. Throughout the explored parameter space, the correlation energy is initially negative and increases during the time evolution, attaining positive average values for some parameter configurations. Figure 2(b) thus visualizes the buildup of two‑particle correlations. The cut‑off at $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$ was again chosen empirically to separate regimes of pronounced correlation‑energy increase from those exhibiting only a modest rise. Both contour lines in Fig. 2 indicate structures in the dynamics of the 2RDM across the parameter plane, which later help to delimit regimes of positive and negative correlations in the dynamics of the cumulants.
@@
Figure 2: We study the six-site Fermi-Hubbard model of Eq. (1). Both plots show a parameter scan over the initial potential strength $V$ and the interaction $U$. The dynamical buildup of three-particle correlations as described in Eq. (10) is visualized in (a). The solid black line corresponds to $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$, which indicates the crossover between moderate and strong correlation buildup. In (b) the time-averaged correlation energy relative to the initial excitation energy $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)$ is shown, where the black dashed line corresponds to $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$. This delimits the regime of strong two-particle correlation buildup. The color-bar is cut below $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$ to have a clear distinction between the two regimes. Both values were chosen in accordance with Fig. 3 since they also delimit the positive and negative as well as the strong and weak correlations of the cumulant dynamics, respectively.
@@
We hypothesize that in regions with high (anti-)correlation between two-particle and three-particle cumulants, the neural ODE should perform well: The information about the three-particle cumulant, which is required to close the equations of motion for the 2RDM, should be contained in the time-local 2RDM. In the correlated regime, this assumption seems natural. In the anti-correlated regime, there is a strong correlation between these two quantities, though negative. This could hint at the existence of a time-local reconstruction functional, which takes this inverse relation into account. The current reconstruction functionals only perform well in the positively correlated regime, since they are proportional to polynomials of the two-particle cumulant. The existence of such an inverse functional would allow for efficient numerical integration of the equations of motion with accurate results even in the regime of large quenches and large correlation buildups. The neural ODE serves as a model-agnostic tool: It can only learn Markovian dynamics by design. Thus, if memory is required to reconstruct the three-particle cumulant and hence close the equations of motion, the model should not be able to learn and predict the exact dynamics of the 2RDM.
@@
To assess this hypothesis, we look at the Pearson correlation coefficient $C_{f,g}$ [74], which provides a quantitative measure of the linear relationship between two time‑dependent observables $f(t)$ and $g(t)$. It is defined as the normalized covariance
@@
By construction, $1\leq C_{f,g}\leq 1$; $C_{f,g}=1$ (1-1) indicates perfect (anti‑)correlation, while $C_{f,g}=0$ corresponds to the absence of correlation. In the present work, $C_{f,g}$ is employed to assess the temporal correlation between the size of the two‑particle cumulant $\Delta_{12}(t)$ and the size of the kernel component of the three‑particle cumulant $\Delta_{123,\text{K}}(t)$ as measured by the Frobenius norm. Since the initial rise of the cumulants is always correlated and we are only interested in the correlations of the steady-state fluctuations, we omit the start of the time series and empirically choose $t_{0}=10\,J^{-1}$. According to our hypothesis, a high absolute value of $C_{\Delta_{12},\Delta_{123,\text{K}}}$ supports the validity of time‑local reconstruction functionals, whereas low absolute values signal the need for memory‑dependent extensions. In Fig. 3(a) and (b) the Pearson correlation coefficient between the Frobenius norms of the kernel component of the three-particle cumulant $\Delta_{123,\text{K}}^{\uparrow\uparrow\downarrow}$ and the two-particle cumulants $\Delta_{12}^{\uparrow\uparrow}$ and $\Delta_{12}^{\uparrow\downarrow}$, respectively, are shown. Both contour lines, which are equal to those in Fig. 2, help to delimit the crossover between positive and negative correlation between the cumulant dynamics. Note that evaluating in Fig. 2 and 3 the entire three-particle cumulant instead of only the kernel component (such as in Ref. [34]) leads to almost the same results because the trace information is a small, practically time-independent contribution.
Figure 3: In (a) and (b) the Pearson correlation coefficient $C_{\Delta_{12},\Delta_{123,\text{K}}}$ between the Frobenius norms of the kernel component of the three-particle cumulant $\Delta_{123,\text{K}}^{\uparrow\uparrow\downarrow}$ and the two-particle cumulants $\Delta_{12}^{\uparrow\uparrow}$ and $\Delta_{12}^{\uparrow\downarrow}$, respectively, are shown across the $U$-$V$ plane for the six-site Fermi-Hubbard model of Eq. (1). The contour lines $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$ and $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$ from Fig. 2 are visualized as well. In (c) and (d) the Pearson correlation coefficient between the predicted and the exact time series of the 2RDM $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ is visualized for a prediction length of $t_{\text{pred}}=20\,J^{-1}$ and $t_{\text{pred}}=25\,J^{-1}$ (2000 and 2500 time steps), respectively, starting from $D_{12}(t=40\,J^{-1})$. For each parameter configuration, one model was trained.
We use $D_{12}(t_{0}=40\,J^{-1})$ as the initial input to the neural ODE, which is then employed to predict the subsequent evolution of the 2RDM starting from $t_{0}=40\,J^{-1}$. Consequently, we define the prediction length as
@@
To quantify the predictive performance of the neural ODE, we calculate the Pearson correlation coefficient between the predicted time series of the 2RDM and the actual time series for the reduced representation (upper triangular values only) $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ for a prediction length of $t_{\text{pred}}=20\,J^{-1}$ and $t_{\text{pred}}=25\,J^{-1}$ (2000 and 2500 time steps) starting from $D_{12}(t=40\,J^{-1})$, visualized in Fig. 3(c) and (d), respectively. The contour lines are the same as in Fig. 2. For a prediction length of $t_{\text{pred}}=20\,J^{-1}$ the performance is high in the correlated regime and reasonably good in the anti-correlated regime. However, for a longer prediction length of $t_{\text{pred}}=25\,J^{-1}$, the performance in the anti-correlated regime deteriorates significantly, while it remains good in the correlated regime. Notably, in the anti-correlated region, values of the Pearson coefficient $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}\in[0.4,0.7]$ are observed. This means that time-local neural ODE still captures a substantial part of the dynamics in these regions, which suggests that the non-Markovian memory effects are weak or short-lived. In such parameter regions, reconstruction functionals that incorporate only a short memory window may already suffice for an accurate closure of the BBGKY hierarchy.
The Pearson correlation coefficient is used only as a post-processing metric to evaluate how accurately the neural ODE reproduces the 2RDM time series. The model is trained by minimizing the mean squared error between predicted and reference trajectories, as defined in Eq. (9). This training loss decreases by several orders of magnitude for the entire parameter scan, such that we regard the optimization as successful for all UVU-V values. Given the successful training, we then use the Pearson coefficient as a diagnostic for effective Markovianity: since neural ODEs learn time-local (Markovian) dynamics, the coefficient should be high when the underlying dynamics is effectively Markovian, and lower when non-Markovian memory effects are relevant.
To further investigate the accuracy of the prediction, we calculate two quantities from the predicted time series, namely the occupation number $n(t)$ and the doublon occupation number $d(t)$ (number of pairs on a site).
Since all sites show similar dynamics, we only consider site 1. The accuracy of the prediction is determined by the time-integrated deviations from the exact result
@@
This quantity as well as the deviations for the doublon occupation number (i.e. occupation of sites by two-particles) $\overline{\delta d_{1}}$ are shown in Fig. 4 for both prediction lengths. Again, we can see that the error is low in the correlated regime, while it is overall higher (and more strongly fluctuating) in the anti-correlated regime. Figure 4 also depicts that even for a shorter prediction length, there is a significant difference in the accuracy of the prediction in the correlated and anti-correlated regimes.
These results contradict our hypothesis and thus suggest that an accurate time-local reconstruction functional cannot be found in the anti-correlated regime. The solid contour line separates moderate from strong correlations buildup of the kernel component of the three-particle cumulant ($\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$). In both Fig. 3 and 4, this line quite accurately localizes regions of good and bad performance, indicating that the existence of a time-local reconstruction functional depends on this quantity.
@@
Figure 5: Line scan at fixed U=1.6JU=1.6\,J, with the Pearson correlation coefficient $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ as a function of $V$. Comparison between the original prediction from Fig. 3 and the prediction obtained by a hyper-parameter-optimized model at each $V$. The hyper-parameter-optimization clearly suppresses fluctuations while preserving the overall trend. The dashed and solid vertical lines correspond to the intersections with the contour lines $\bar{E}_{\text{corr}}/E_{\text{pot}}(0)=-0.1$ and $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}=0.65$, respectively, from Fig. 2.
To investigate whether hyper-parameter optimization reduces the fluctuations, we fixed U=1.6JU=1.6\,J and performed a line scan over $V$. For each value of $V$, we performed hyper-parameter optimization. Figure 5 compares the original result from Fig. 3 with the optimized result. The optimization substantially reduces fluctuations and supports the robustness of the trends discussed above. Since the Pearson correlation coefficient slightly decreases towards the dashed line, similar to the correlation between the cumulants in Fig. 3(b), this suggests that the dynamics is Markovian to a good approximation in the correlated regime and memory is required as soon as this correlation decreases. For longer prediction horizons even small deviations from Markovianity can accumulate, which explains the shift between Fig. 5 (a) and (b). It is remarkable that the hyper-parameter optimization eliminates the noise in the prediction, making the tuned model less uncertain.
The training of the neural ODE based on the full data as performed in this study is computationally costly. While it is not within the scope of the present study to focus on the computational cost, there are some obvious strategies for cost reduction: (i) reducing the dimensionality of the input data and (ii) reducing the model size. However, we intentionally do not prioritize these strategies in the present work: dimensionality reduction may discard information that is crucial for assessing Markovianity, and therefore could bias the main diagnostic objective. Likewise, substantially smaller networks introduce a stronger information bottleneck, which is not desired in our setup because it can limit representational capacity and confound the interpretation of memory effects.